At 300 K the vapour pressure of a pure liquid A is 553 Torr, of pure liquid B it is 380 Torr. A and B form ideal liquid mixtures.

a) Calculate the vapour pressures of A and B and the total pressure above liquid mixtures which are 25, 50, and 75 mol% B.

b) For a particular liquid mixture, the vapour composition is found to be 0.35 mol% A. Calculate the liquid composition in equilibrium with vapour of this composition. What do you notice about the liquid composition (relative to that of the vapour)?
Answer #1

a)

vapour pressure of pure liquid A = 553 Torr

vapour pressure of pure liquid B = 380 Torr

1) 25 mol% B

mol% = (mole fraction)\times100

mole fraction = \frac{mol percentage}{100}

mole fraction of B (\chi _{B}) = \frac{mol percentage of B}{100}

=  \frac{25}{100}

= 0.25

mole fraction of A (\chi _{A}) = 1- mole fraction B (since there are only two components)

= 1- 0.25

= 0.75

vapour pressure of A in solution = (\chi _{A})\times(vapour pressure of pure liquid A)

= (0.75)\times( 553 Torr)

= 414.75 Torr

vapour pressure of B in solution = (\chi _{B})\times(vapour pressure of pure liquid B)

= (0.25)\times( 380 Torr)

= 96 Torr

vapour pressure of solution = (\chi _{A})\times(vapour pressure of pure liquid A)+(\chi _{B})\times(vapour pressure of pure liquid B)

= (0.75)\times( 553 Torr)+ (0.25)\times( 380 Torr)

= 414.75 Torr +96 Torr

= 510.75 Torr

2) 50 mol% B

mol% = (mole fraction)\times100

mole fraction = \frac{mol percentage}{100}

mole fraction of B (\chi _{B}) = \frac{mol percentage of B}{100}

=  \frac{50}{100}

= 0.50

mole fraction of A (\chi _{A}) = 1- mole fraction B (since there are only two components)

= 1- 0.50

= 0.50

vapour pressure of A in solution = (\chi _{A})\times(vapour pressure of pure liquid A)

= (0.50)\times( 553 Torr)

= 276.5 Torr

vapour pressure of B in solution = (\chi _{B})\times(vapour pressure of pure liquid B)

= (0.50)\times( 380 Torr)

= 190 Torr

vapour pressure of solution = (\chi _{A})\times(vapour pressure of pure liquid A)+(\chi _{B})\times(vapour pressure of pure liquid B)

= (0.50)\times( 553 Torr)+ (0.50)\times( 380 Torr)

= 276.5 Torr +190 Torr

= 466.5 Torr

3) 75 mol% B

mol% = (mole fraction)\times100

mole fraction = \frac{mol percentage}{100}

mole fraction of B (\chi _{B}) = \frac{mol percentage of B}{100}

=  \frac{75}{100}

= 0.75

mole fraction of A (\chi _{A}) = 1- mole fraction B (since there are only two components)

= 1- 0.75

= 0.25

vapour pressure of A in solution = (\chi _{A})\times(vapour pressure of pure liquid A)

= (0.25)\times( 553 Torr)

= 133.25 Torr

vapour pressure of B in solution = (\chi _{B})\times(vapour pressure of pure liquid B)

= (0.75)\times( 380 Torr)

= 285 Torr

vapour pressure of solution = (\chi _{A})\times(vapour pressure of pure liquid A)+(\chi _{B})\times(vapour pressure of pure liquid B)

= (0.25)\times( 553 Torr)+ (0.75)\times( 380 Torr)

= 133.25 Torr +285 Torr

= 418.25 Torr

b)

Let total vapour pressure = Pt

pressure due to A vapour = PA

pressure due to B vapour = PB

let mole fraction of A vapour = \phi _{A} = 0.35 % = 0.0035

let mole fraction of B vapour = \phi _{B} = 1-\phi _{A} = 1- 0.0035 = 0.9965

vapour pressure due to A = (\chi _{A})\times(vapour pressure of pure liquid A)

= (\chi _{A})\times( 553 Torr)

since T and V are constant

pressure due to A vapour =  vapour pressure due to A = (mole fraction of A vapour)\times(total vapour pressure)

= \phi _{A}\times Pt

= 0.0035\times Pt

form above equations

(\chi _{A})\times( 553 Torr) = 0.0035\times Pt

  Pt = \frac{\chi _{A}\times 553}{0.0035}

= 1.58 \times 105\chi _{A} (1)

vapour pressure due to B = (\chi _{B})\times(vapour pressure of pure liquid A)

= (\chi _{B})\times(380)

pressure due to A vapour =  vapour pressure due to A = (mole fraction of B vapour)\times(total vapour pressure)

= \phi _{B}\times Pt

= 0.9965\times Pt

therefore   (\chi _{B})\times(380) = 0.9965\times Pt

Pt = \frac{\chi _{B}\times 380}{0.9965}

= 381,33 \chi _{B} (2)

from 1 and 2   1.58 \times 105\chi _{A} = 381.33 \chi _{B}

\chi _{A} = 414.339 \chi _{B}

but since \chi _{A} and \chi _{B} are mole fractions

\chi _{A}+\chi _{B} = 1

substituting \chi _{A} = 414.339 \chi _{B} in above equation

  414.339 \chi _{B}  +\chi _{B} = 1

415.339 \chi _{B} = 1

\chi _{B} = \frac{1}{415.339}

\chi _{B} = 0.0024

we know that   \chi _{A}+\chi _{B} = 1

  \chi _{B} = 1- \chi _{A}

= 1- 0.0024

= 0.9976

  mole fraction of A (\chi _{A}) = 0.0024

  mole fraction of B (\chi _{B}) = 0.9976